Use the distributive property to rewrite each of the following expressions in expanded form.
a. 8x(x+9)=
b. 7y(3x-2y)=
c. a(a^2b+b^3c)

Answer by Theo(13342)

a. 8x(x+9) = (8x * x) + (8x * 9) = 8x^2 + 72x

b. 7y(3x-2y) = (7y*3x) - (7y*2y) = 21xy - 14y^2


c. a(a^2b+b^3c) = (a*a^2b) + (a*b^3c) = a^3b + ab^3c